OpenStudy (anonymous):

cos150degrees - 3 tan300 degrees + 2 sin 90degrees any mathematician can help me here

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

150 degree comes in second quadrent

OpenStudy (anonymous):

i found the value already but how can i sovle this....wait i will type

OpenStudy (gorv):

so cos give negative value in second quad

OpenStudy (anonymous):

\[\frac{ -\sqrt{3} }{ 2 }- 3(-\sqrt{3}) + 2\]

OpenStudy (anonymous):

i found this value but how can i simplify this bro..do you know please help me buddy :D

OpenStudy (gorv):

yeah ur right

OpenStudy (gorv):

okk i will

OpenStudy (gorv):

take 3 inside the bracket

OpenStudy (gorv):

both term have minus sign and are in multiplication

OpenStudy (gorv):

whenever there are 2 term and both are in multiplication then minus sign become plus

OpenStudy (anonymous):

can u draw or write it so i can understand it quikly buddy :D

OpenStudy (gorv):

okk lets replace sqrt by x

OpenStudy (gorv):

\[\frac{ -x }{ 2 }-3*(-x)+2\]

OpenStudy (gorv):

\[\frac{ -x }{ 2 }+3x+2\]

OpenStudy (gorv):

can u do it now ??

OpenStudy (anonymous):

so that means \[\frac{ -\sqrt{3} }{ 2 }+3\sqrt{3} + 2\]

OpenStudy (anonymous):

is that the lowest tem for it?

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

\[\sqrt{3}*(\frac{ -1 }{ 2 }+3)+2\]

OpenStudy (gorv):

\[\frac{ 5 }{ 2 }*\sqrt{3}+2=\frac{ 5\sqrt{3} +4}{ 2 }\]